### v1

@@ Abstract
Given an Abelian group $\mathcal{G}$, a Boolean-valued function $f:\mathcal{G}\to\{-1,+1\}$, is said to be $s$-sparse, if it has at most $s$-many non-zero Fourier coefficients over the domain $\mathcal{G}$. In a seminal paper, Gopalan et al. [15] proved “Granularity” for Fourier coefficients of Boolean valued functions over $\mathbb{Z}_{2}^{n}$, that have found many diverse applications in theoretical computer science and combinatorics. They also studied structural results for Boolean functions over $\mathbb{Z}_{2}^{n}$ which are approximately Fourier-sparse. In this work, we obtain structural results for approximately Fourier-sparse Boolean valued functions over Abelian groups $\mathcal{G}$ of the form, $\mathcal{G}:=\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}$, for distinct primes $p_{i}$. We also obtain a lower bound of the form $1/(m^{2}s)^{\lceil\varphi(m)/2\rceil}$, on the absolute value of the smallest non-zero Fourier coefficient of an $s$-sparse function, where $m=p_{1}\cdots p_{t}$, and $\varphi(m)=(p_{1}-1)\cdots(p_{t}-1)$. We carefully apply probabilistic techniques from [15], to obtain our structural results, and use some non-trivial results from algebraic number theory to get the lower bound.
We construct a family of at most $s$-sparse Boolean functions over $\mathbb{Z}_{p}^{n}$, where $p>2$, for arbitrarily large enough s, where the minimum non-zero Fourier coefficient is $o(1/s)$. The “Granularity” result of Gopalan et al. implies that the absolute values of non-zero Fourier coefficients of any $s$-sparse Boolean valued function over $\mathbb{Z}_{2}^{n}$ are $\Omega\left({1}/{s}\right)$. So, our result shows that one cannot expect such a lower bound for general Abelian groups.
Using our new structural results on the Fourier coefficients of sparse functions, we design an efficient sparsity testing algorithm for Boolean function, which tests whether the given function is $s$-sparse, or $\epsilon$-far from any sparse Boolean function, and it requires $\mathsf{poly}((ms)^{\varphi(m)},1/\epsilon)$-many queries. Further, we generalize the notion of degree of a Boolean function over the Abelian group $\mathcal{G}$. We use it to prove an $\Omega(\sqrt{s})$ lower bound on the query complexity of any adaptive sparsity testing algorithm.
keywords:
Fourier coefficients, sparse, Abelian, granularity
@@ 1 Introduction
The case of Boolean function complexity with $\mathcal{D}=\mathbb{Z}_{2}^{n}$ has been widely studied. These functions are often analyzed in connection with their Fourier transform (see the archived version of the paper, cite) and a significant amount of research has focused on the structural properties of Fourier spectra of important classes of these functions. One such class that this work focuses on is that of Fourier-sparse functions. These are functions with only a few non-zero Fourier coefficients, formally defined in the archived version of the paper, cite. We will denote by $s_{f}$ the Fourier sparsity of a Boolean function $f$. Fourier sparsity and Fourier-sparse functions have known connections with a variety of areas of Boolean function analysis and computational complexity like property testing [15], learning theory [21, 3], distance estimation [38] and communication complexity [25, 34, 20, 29, 24]. These connections provide enough motivation to comprehend the structure of Fourier coefficients for Fourier-sparse Boolean functions.
@@
In this work, we extend the study of Fourier-sparse Boolean functions to the domains $\mathcal{D}$, which are Abelian groups of the form $\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}$, where $p_{1},\ldots,p_{t}$ are distinct prime numbers. Boolean functions over general Abelian groups have been studied in both mathematics and computer science. A celebrated result regarding such functions is Chang’s Lemma [10]. Chang’s lemma over $\mathbb{Z}_{2}^{n}$ has found numerous applications in complexity theory and algorithms [6, 9], analysis of Boolean functions [17, 34], communication complexity [34, 22], extremal combinatorics [13], and many more. Recently, [8] improved Chang’s lemma over $\mathbb{Z}_{2}^{n}$ for some special settings of parameters, where Fourier sparsity played a crucial role. One motivation to study Fourier sparsity over a broader class of Abelian groups is to investigate possible generalizations of their bounds to those groups.
@@
Fourier analysis over finite Abelian groups in cryptography. For the past three decades, the field of cryptography has been utilizing concepts derived from Fourier analysis, specifically over finite Abelian groups. Akavia, Goldwasser, and Safra [2] have combined some of these concepts to develop a comprehensive algorithm that can detect “large” Fourier coefficients of any concentrated function on finite Abelian groups, and compute a sparse approximation for the same. This algorithm has gained significant attention within the cryptography community, especially regarding the notion of “bit security” of the discrete logarithm problem (DLP), RSA, and learning with errors (LWE) problems; see [27, 14, 1]. In particular, the “nice” structural results on the Fourier coefficients of a Boolean-valued function over the general Abelian group are of utmost importance and interest from a crypto-theoretic point of view.
@@
Interestingly, there are strong relationships between learning, sparsity, and sampling, in the context of Fourier-sparse Boolean functions. They have been rigorously studied in [31, 32, 33]. In [31], the authors asked the following:
Question 1.1.
@@
What can be said about the structure of the Fourier coefficients of a Boolean function $f$ over an Abelian group $\mathcal{G}=\mathbb{Z}_{p}^{n}$, where the support is significantly smaller compared to $\mathcal{G}$?
@@
Gopalan et al. [15] proved that any non-zero Fourier coefficient of a Boolean function over $\mathbb{Z}_{2}^{n}$ with Fourier sparsity at most $s_{f}$, is at least $\frac{1}{s_{f}}$ in its absolute value. This gave a satisfactory answer of 1.1 over $\mathbb{Z}_{2}^{n}$. Furthermore, they proved robust versions of their result for functions which are approximately Fourier-sparse. Finally, their structural results were used to design a sample-efficient algorithm to test whether a function is Fourier-sparse.
@@
In our work, we undertake the same task for Boolean functions over Abelian groups of the form $\mathbb{Z}_{p}^{n}$, for a prime $p\geq 3$. We prove lower bounds on the absolute value of any non-zero coefficient in terms of $s_{f}$ and $p$. We also prove a tightness result that complements our lower bounds. In particular, our bound implies that a lower bound of $\frac{1}{\Theta(s_{f})}$ that [15] showed does not hold anymore for $\mathbb{Z}_{p}^{n}$. Finally, we use our bounds to design a testing algorithm for Fourier-sparse Boolean functions over $\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}$. This is probably the first time, we have advanced on understanding some structure on the Fourier-sparse coefficients over a general Abelian group, and shed some light on 1.1.
It is well-known that any Abelian group is isomorphic to a group of the form $\mathbb{Z}_{p_{1}^{n_{1}}}\times\cdots\times\mathbb{Z}_{p_{t}^{n_{t}}}$ where $p_{1},\ldots,p_{t}$ are prime numbers and $n_{1},\ldots,n_{t}$ are positive integers. Unfortunately, our techniques fall short of handling even a simpler case of $\mathbb{Z}_{p^{2}}$, fundamentally because of the absence of a linear structure on the solution space of systems of equations over $\mathbb{Z}_{p^{2}}$. Therefore, we are leaving the task of investigating the existence of similar bounds and algorithmic results when the domain of the function is a general Abelian group to future research.
@@
There has been a considerable amount of interest in studying the complexity of reconstructing or learning functions of the form $f:\mathcal{D}\to\mathbb{C}$, where $\mathcal{D}$ is a known domain (more general than a hypercube, such as a general finite Abelian group), and $f$ is Fourier-sparse; see [33, 28, 26, 11, 37]. Fourier-sparse functions over various finite Abelian groups have gained much interest with the advancement in sparse Fourier transform algorithms [18, 19, 5]. These algorithms improve the efficiency of the standard Fast Fourier Transform algorithms by taking advantage of the sparsity itself. To reliably use sparse Fourier transform algorithms, it is beneficial to have a way to test if a function is $s$-sparse or, more generally, to estimate the distance of a function to the closest $s$-sparse function. In this work, we consider the problem of non-tolerant sparsity-testing of Boolean functions over finite Abelian groups.
Finally, apart from mathematical curiosity and potential cryptographic applications (as mentioned previously), structural results on Fourier-sparse functions $f:\mathbb{Z}_{N}\to\mathbb{C}$, for some $N\in\mathbb{N}$, have also found algorithmic applications in SOS-optimization and control theory. These applications have further implications in certifying maximum satisfiability (MAX-SAT) and maximum k-colorable subgraph (MkCS) problems; see [12, 36, 35].
@@ 1.1 Our results
Throughout the article, we will be working with the Abelian group $\mathcal{G}:=\mathbb{Z}_{p_{1}}^{n_{1}}\times\dots\times\mathbb{Z}_{p_{t}}^{n_{t}}$ where $p_{i}$ are primes. Let $f:\mathcal{G}\to\{-1,+1\}$, and $f(x)=\sum_{\chi\in\hat{\mathcal{G}}}\widehat{f}(\chi)\chi(x)$ be the Fourier transform of $f$, where $\hat{\mathcal{G}}$ is the set of characters of the Abelian group $\mathcal{G}$.
We say a Boolean function is $s$-sparse, if it has at most $s$ non-zero Fourier coefficients. [15] proved that for any $s$-sparse Boolean functions over $\mathbb{Z}_{2}^{n}$, the magnitude of the Fourier coefficients are $k$-granular where $k=\lceil\log_{2}s\rceil+1$. A real number is $k$-granular if it is an integer multiple of $1/2^{k}$. One wonders whether such a phenomenon still holds over a more general group $\mathcal{G}$.
@@
Definition 1.1.
(Granularity) A complex number is said to be $k$-granular or has granularity $k$ with respect to $\mathbb{Z}_{p}$ if it is of the form $\frac{g(\omega_{p})}{p^{k}}$, where $g(X)\in\mathbb{Z}[X]$ and $\omega_{p}$ is a primitive $p^{th}$ root of unity.
More generally, a complex number is said to be $(m_{1},\ldots,m_{t})$-granular with respect to $(\mathbb{Z}_{p_{1}},\dots,\mathbb{Z}_{p_{t}})$ if it is of the form $\frac{g(\omega_{p_{1}},\ldots,\omega_{p_{t}})}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}$, where $g(X_{1},\ldots,X_{t})\in\mathbb{Z}[X_{1},\ldots,X_{t}]$ and $\omega_{p_{i}}$ is a primitive $p_{i}^{th}$ root of unity, $i\in[k]$.
Note that, this goes well with the definition of granularity in [15] for the case of $\mathbb{Z}_{2}$ as $\omega_{2}$ is either $+1$ or $-1$ and hence $g(\omega_{2})$ is an integer for any $g(X)\in\mathbb{Z}[X]$.
A function $f:\mathcal{G}\to\{-1,+1\}$ is said to be $(m_{1},\dots,m_{t})$-granular with respect to $(\mathbb{Z}_{p_{1}},\ldots,\mathbb{Z}_{p_{t}})$ if each Fourier coefficient of $f$ is $(m_{1},\dots,m_{t})$-granular.
@@
($\mu$-close to $k$-granular) A complex number $v$ is said to be $\mu$-close to $k$-granular with respect to $\mathbb{Z}_{p}$ if $|v-\frac{g(\omega_{p})}{p^{k}}|\leq\mu$, for some non-zero polynomial $g(X)\in\mathbb{Z}[X]$.
Note that a similar notion can be defined for the case of $\mu$-close to $(m_{1},\ldots,m_{t})$-granular with respect to $(\mathbb{Z}_{p_{1}},\dots,\mathbb{Z}_{p_{t}})$.
Now, we are ready to formally state our two main structural results. All our results hold for Boolean-valued functions over the more general Abelian group $\mathcal{G}$. But for simplicity of presentation and ease of understanding, we first present the result for Boolean-valued functions over $\mathbb{Z}_{p}^{n}$. Later, in Appendix F, we present more general versions of our results.
Our first theorem says that for any Boolean-valued function over $\mathbb{Z}_{p}^{n}$ that is close to being sparse, all its large Fourier coefficients are close to being granular. This is a generalization of the structural theorem of [15, Theorem 3.3], which was proved over $\mathbb{Z}_{2}^{n}$.
@@
Let $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ be a Boolean-valued function and let $B$ be the set of characters corresponding to the set of $s$-largest Fourier coefficients of $f$ (in terms of magnitude). If $\sum_{\chi\in B}|\widehat{f}(\chi)|^{2}\geq(1-\mu)$ then for all $\chi\in B$, $\widehat{f}(\chi)$ is $\frac{\mu}{\sqrt{s}}$-close to $\lceil\log_{p}s\rceil+1$-granular.
For a function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, $\sum_{\chi\in B}|\widehat{f}(\chi)|^{2}\geq(1-\mu)$ (where $B$ is the set of characters corresponding to the set of $s$ largest coefficients of $f$) can also be stated as “there is an $s$-sparse function $g:\mathbb{Z}_{p}^{n}\to\mathbb{C}$ with the $\ell_{2}$-distance between $f$ and $g$ is at most $\sqrt{\mu}$”. But note that this does not guarantee that there is an $s$-sparse Boolean-valued function $g:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ with $\ell_{2}$-distance between $f$ and $g$ being at most $\sqrt{\mu}$. However, our second theorem proves that one can indeed find an $s$-sparse Boolean-valued function in a close enough vicinity, thus generalizing [15, Theorem 3.4].
@@
Let $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ be a Boolean-valued function and let $B$ be the set of characters corresponding to the set of $s$-largest Fourier coefficients of $f$ (in terms of magnitude). If $\sum_{\chi\in B}|\widehat{f}(\chi)|^{2}\geq(1-\mu)$, with $\mu\leq\frac{1}{8(p^{2}s)^{p-1}}$ then there exists an $s$-sparse Boolean-valued function $F:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ with $\ell_{2}$ distance between $f$ and $F$ is at most $\sqrt{2}\mu$.
Theorem 1.3 and Theorem 1.4 can be suitably generalized to functions from $\mathcal{G}$ to $\{-1,+1\}$. But for the clarity of presentation, we state the generalized theorems and present their proofs in Appendix F.
One important corollary of Theorem 1.3 is that for any $s$-sparse Boolean function $f:\mathbb{Z}_{2}^{n}\to\{-1,+1\}$ the non-zero Fourier coefficients has magnitude at least $1/2^{k}$, where $k=\lceil\log_{2}s\rceil+1$. Unfortunately, such a simple corollary cannot be claimed for $s$-sparse functions $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ if $p\neq 2$. The main reason is that the definition of granularity is for complex numbers, rather than real numbers and hence such a lower bound cannot be directly deduced. However, borrowing results from algebraic number theory, we can obtain a lower bound on the Fourier coefficients of Boolean-valued functions from $\mathbb{Z}_{p}^{n}$ to $\{-1,+1\}$ for any arbitrary prime $p$.
@@
Let $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, with Fourier sparsity $s_{f}$. Then, for any $\chi\in\operatorname{supp}(f)$, we have $|\hat{f}(\chi)|\geq\frac{1}{(p^{2}s_{f})^{\lceil(p-1)/2\rceil}}$
{remark*}
One can also prove a weaker lower bound of the form $1/((s_{f}+1)\sqrt{s_{f}})^{s_{f}}$, which is $p$-independent; for details, see Theorem C.8. Observe that the lower bound in Theorem 1.5 is much lower than $1/s_{f}$. One may wonder how tight our result is. It is known that $1/s$ is tight for the case when $p=2$. For example, consider the function $AND:\mathbb{Z}_{2}^{n}\to\{-1,+1\}$. Its non-empty Fourier coefficients are either $\frac{1}{2^{n-1}}$ or $-\frac{1}{2^{n-1}}$, while the empty (constant) coefficient being $1-\frac{1}{2^{n-1}}$.
@@
We prove a generalized version of the lower bound result (Theorem 1.5) for Boolean-valued functions over $\mathcal{G}$. The above example (from Theorem 1.6) can also be easily extended to obtain Boolean-value functions over $\mathcal{G}$ demonstrating similar bounds on the absolute value of the non-trivial Fourier coefficients.
Finally, we design efficient algorithms for testing whether a function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ is $s$-sparse or “far” from $s$-sparse-Boolean. To state our results we need to define what we mean by a function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ is $\epsilon$-far from $s$-sparse.
@@
A function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ is $\epsilon$-far from $s$-sparse-Boolean if for every $s$-sparse function $g:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ the $\ell_{2}$-distance of $f$ and $g$ is at least $\sqrt{\epsilon}$. [^1] In property testing usually the distance measure used is Hamming distance between two Boolean functions. But since we are using $\ell_{2}$ distance in our other theorem, so for ease of presentation, we have defined the farness in terms of $\ell_{2}$ instead of Hamming distance. Also, note that for a pair of Boolean-valued functions the square of the $\ell_{2}$ distance and Hamming distance are the same up to a multiplicative factor of $4$, see Lemma A.15.
We say that an algorithm (property tester) $\mathcal{A}$  $\epsilon$-tests $\mathcal{C}$, for a class of functions $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, if given access to the truth table of a function $f$, whether $f\in\mathcal{C}$, or $f$, is “$\epsilon$-far from $\mathcal{C}$” can be tested using $\mathcal{A}$ with success probability (called the confidence) $\geq 2/3$. The number of queries to the truth-table of $f$ made by $\mathcal{A}$ is called the query complexity of $\mathcal{A}$.
Using the structure theorems (Theorem 1.3 and Theorem 1.4), we prove the following theorem which tests sparsity of a Boolean-valued function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$.

### v2

@@ Abstract
We study Fourier-sparse Boolean functions over general finite Abelian groups. A Boolean function $f:\mathcal{G}\to\{-1,+1\}$ is $s$-sparse if it has at most $s$ non-zero Fourier coefficients. We introduce a general notion of granularity of Fourier coefficients and prove that every non-zero coefficient of an $s$-sparse Boolean function has magnitude at least
$\frac{1}{2^{\varphi(\Delta)/2}\,s^{\varphi(\Delta)/2}},$
where $\Delta$ denotes the exponent of the group $\mathcal{G}$ (that is, the maximum order of an element in $\mathcal{G}$) and $\varphi$ is the Euler’s totient function. This generalizes the celebrated result of Gopalan et al. (SICOMP 2011) for $\mathbb{Z}_{2}^{n}$, extending it to all finite Abelian groups via new techniques from group theory and algebraic number theory.
Using our new structural results on the Fourier coefficients of sparse functions, we design an efficient sparsity testing algorithm for Boolean functions. The tester distinguishes whether a given function is $s$-sparse or $\epsilon$-far from every $s$-sparse Boolean function, with query complexity $\mathsf{poly}\left((2s)^{\varphi(\Delta)},1/\epsilon\right)$. In addition, we generalize the classical notion of Boolean degree to arbitrary Abelian groups and establish an $\Omega(\sqrt{s})$ lower bound for adaptive sparsity testing.
@@ 1 Introduction
The case of Boolean function complexity with $\mathcal{D}=\mathbb{Z}_{2}^{n}$ has been widely studied. These functions are often analyzed in connection with their Fourier transform (see Section 2.1) and a significant amount of research has focused on the structural properties of Fourier spectra of important classes of these functions. One such class that this work focuses on is that of Fourier-sparse functions. These are functions with only a few non-zero Fourier coefficients, formally defined in Definition 2.11. We will denote by $s_{f}$ the Fourier sparsity of a Boolean function $f$. Fourier sparsity and Fourier-sparse functions have known connections with a variety of areas of Boolean function analysis and computational complexity like property testing [GOS+11], learning theory [HR16, ACL+21a], distance estimation [YZ20] and communication complexity [MO09, TWXZ13, HHL18, San19, MS20]. These connections provide enough motivation to comprehend the structure of Fourier coefficients for Fourier-sparse Boolean functions.
@@
In this work, we extend the study of Fourier-sparse Boolean functions to the domains $\mathcal{D}$, which are finite Abelian groups. Boolean functions over general Abelian groups have been studied in both mathematics and computer science. A celebrated result regarding such functions is Chang’s Lemma [Cha02]. Chang’s lemma over $\mathbb{Z}_{2}^{n}$ has found numerous applications in complexity theory and algorithms [BRTW14, CLRS16], analysis of Boolean functions [GS08, TWXZ13], communication complexity [TWXZ13, HLY19], extremal combinatorics [FKKK18], and many more. Recently, [CMM+21] improved Chang’s lemma over $\mathbb{Z}_{2}^{n}$ for some special settings of parameters, where Fourier sparsity played a crucial role. One motivation to study Fourier sparsity over a broader class of Abelian groups is to investigate possible generalizations of their bounds to those groups.
@@
Fourier analysis over finite Abelian groups in cryptography. For the past three decades, the field of cryptography has been utilizing concepts derived from Fourier analysis, specifically over finite Abelian groups. Akavia, Goldwasser, and Safra [AGS03] have combined some of these concepts to develop a comprehensive algorithm that can detect “large” Fourier coefficients of any concentrated function on finite Abelian groups, and compute a sparse approximation for the same. This algorithm has gained significant attention within the cryptography community, especially regarding the notion of “bit security” of the discrete logarithm problem (DLP), RSA, and learning with errors (LWE) problems; see [Reg09, GLS18, Aka08]. In particular, the “nice” structural results on the Fourier coefficients of a Boolean-valued function over the general Abelian group are of utmost importance and interest from a crypto-theoretic point of view.
@@
Interestingly, there are strong relationships between learning, sparsity, and sampling, in the context of Fourier-sparse Boolean functions. They have been rigorously studied in [Sit94, SS00, SS01]. In [Sit94], the authors asked the following:
Question 1.
@@
What can be said about the structure of the Fourier coefficients of a Boolean function $f$ over a finite Abelian group $\mathcal{G}$, where the support is significantly smaller compared to $\mathcal{G}$?
@@
Gopalan et al. [GOS+11] proved that any non-zero Fourier coefficient of a Boolean function over $\mathbb{Z}_{2}^{n}$ with Fourier sparsity at most $s_{f}$, is at least $\frac{1}{s_{f}}$ in its absolute value. This gave a satisfactory answer of 1 over $\mathbb{Z}_{2}^{n}$. Furthermore, they proved robust versions of their result for functions which are approximately Fourier-sparse. Finally, their structural results were used to design a sample-efficient algorithm to test whether a function is Fourier-sparse.
@@
In our work, we undertake the same task for Boolean functions over finite Abelian groups $\mathcal{G}$. Any finite Abelian group $\mathcal{G}$ can be written in the form $\mathbb{Z}_{{p_{1}}^{m_{1}}}\times\dots\times\mathbb{Z}_{{p_{T}}^{m_{T}}}$, ($p_{i},\ i\in[T]$ are primes and not necessarily distinct). We prove lower bounds on the absolute value of any non-zero coefficient in terms of $s_{f}$ and $\Delta$, where $\Delta=\operatorname{LCM}\{p_{1}^{m_{1}},\ldots,p_{T}^{m_{T}}\}$. Chakraborty et al. [CDD+24] have already proven these results for Abelian groups of the form $\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{T}}^{n_{T}}$, where $p_{i},\ i\in[T]$ are distinct primes, which is a generalization of Gopalan et al. [GOS+11]. Observe that $\mathbb{Z}_{p}^{n}$ for $p$ prime is a vector space. An important thing to note here is that a finite Abelian group is not a vector space in general. Since most of the properties of vector space are lost in case of $\mathcal{G}$, we use different techniques to prove our results. They also showed a tightness result that complements our lower bounds. In particular, our bound implies that a lower bound of $\frac{1}{\Theta(s_{f})}$ that [GOS+11] showed does not hold anymore for $\mathcal{G}$. Finally, we use our bounds to design a testing algorithm for Fourier-sparse Boolean functions over $\mathcal{G}$.
@@
There has been a considerable amount of interest in studying the complexity of reconstructing or learning functions of the form $f:\mathcal{D}\to\mathbb{C}$, where $\mathcal{D}$ is a known domain (more general than a hypercube, such as a general finite Abelian group), and $f$ is Fourier-sparse; see [SS01, RV08, Mor13, CD20, YYZ23]. Fourier-sparse functions over various finite Abelian groups have gained much interest with the advancement in sparse Fourier transform algorithms [HIKP12a, HIKP12b, BMV18]. These algorithms improve the efficiency of the standard Fast Fourier Transform algorithms by taking advantage of the sparsity itself. To reliably use sparse Fourier transform algorithms, it is beneficial to have a way to test if a function is $s$-sparse or, more generally, to estimate the distance of a function to the closest $s$-sparse function. In this work, we consider the problem of non-tolerant sparsity-testing of Boolean functions over finite Abelian groups.
Finally, apart from mathematical curiosity and potential cryptographic applications (as mentioned previously), structural results on Fourier-sparse functions $f:\mathbb{Z}_{N}\to\mathbb{C}$, for some $N\in\mathbb{N}$, have also found algorithmic applications in SOS-optimization and control theory. These applications have further implications in certifying maximum satisfiability (MAX-SAT) and maximum k-colorable subgraph (MkCS) problems; see [FSP15, YYZ22b, YYZ22a].
@@ 1.1 Our results
Throughout the article, we will be working with finite Abelian groups. Any finite Abelian group $\mathcal{G}$ can be written in the form $\mathcal{G}:=\mathbb{Z}_{{p_{1}}^{m_{1}}}\times\dots\times\mathbb{Z}_{{p_{T}}^{m_{T}}}$, where $p_{i}$ are primes and not necessarily distinct. Let $f:\mathcal{G}\to\{-1,+1\}$, and $f(x)=\sum_{\chi\in\hat{\mathcal{G}}}\widehat{f}(\chi)\chi(x)$ be the Fourier transform of $f$, where $\hat{\mathcal{G}}$ is the set of characters of the Abelian group $\mathcal{G}$. Also, let $\Delta=\operatorname{LCM}\{p_{1}^{m_{1}},\ldots,p_{T}^{m_{T}}\}$.
We say a Boolean function is $s$-sparse, if it has at most $s$ non-zero Fourier coefficients. [GOS+11] proved that for any $s$-sparse Boolean functions over $\mathbb{Z}_{2}^{n}$, the magnitude of the Fourier coefficients are $k$-granular where $k=\lceil\log_{2}s\rceil+1$. A real number is $k$-granular if it is an integer multiple of $1/2^{k}$. One wonders whether such a phenomenon still holds over a more general group $\mathcal{G}$.
@@
Definition $1.$
(Granularity) A complex number is said to be $k$-granular or has granularity $k$ with respect to $\mathcal{G}$ if it is of the form $\frac{g(\omega_{\Delta})}{\Delta^{k}}$, where $g(X)\in\mathbb{Z}[X]$ and $\omega_{\Delta}$ is a primitive $(\Delta)^{th}$ root of unity.
Note that, this goes well with the definition of granularity in [GOS+11] for the case of $\mathbb{Z}_{2}$ as $\omega_{2}$ is either +1+1 or 1-1 and hence $g(\omega_{2})$ is an integer for any $g(X)\in\mathbb{Z}[X]$.
@@
($\mu$-close to $k$-granular) A complex number $v$ is said to be $\mu$-close to $k$-granular with respect to $\mathcal{G}$ if $|v-\frac{g(\omega_{\Delta})}{\Delta^{k}}|\leq\mu$, for some non-zero polynomial $g(X)\in\mathbb{Z}[X]$.
Now, we are ready to formally state our two main structural results. All our results hold for Boolean-valued functions over the more general Abelian group $\mathcal{G}$.
Our first theorem says that for any Boolean-valued function over $\mathcal{G}$ that is close to being sparse, all its large Fourier coefficients are close to being granular. This is a generalization of the structural theorem of [GOS+11, Theorem 3.3], which was proved over $\mathbb{Z}_{2}^{n}$.
@@
Let $f:\mathcal{G}\to\{-1,+1\}$ be a Boolean-valued function and let $B$ be the set of characters corresponding to the set of $s$-largest Fourier coefficients of $f$ (in terms of magnitude). If $\sum_{\chi\in B}|\widehat{f}(\chi)|^{2}\geq(1-\mu)$ then for all $\chi\in B$, $\widehat{f}(\chi)$ is $\frac{\mu}{\sqrt{s}}$-close to $k$-granular, where $k=\lceil\log_{\Delta}2s\rceil$.
For a function $f:\mathcal{G}\to\{-1,+1\}$, $\sum_{\chi\in B}|\widehat{f}(\chi)|^{2}\geq(1-\mu)$ (where $B$ is the set of characters corresponding to the set of $s$ largest coefficients of $f$) can also be stated as “there is an $s$-sparse function $g:\mathcal{G}\to\mathbb{C}$ with the $\ell_{2}$-distance between $f$ and $g$ is at most $\sqrt{\mu}$”. But note that this does not guarantee that there is an $s$-sparse Boolean-valued function $g:\mathcal{G}\to\{-1,+1\}$ with $\ell_{2}$-distance between $f$ and $g$ being at most $\sqrt{\mu}$. However, our second theorem proves that one can indeed find an $s$-sparse Boolean-valued function in a close enough vicinity, thus generalizing [GOS+11, Theorem 3.4].
@@
Let $f:\mathcal{G}\to\{-1,+1\}$ be a Boolean-valued function and let $B$ be the set of characters corresponding to the set of $s$-largest Fourier coefficients of $f$ (in terms of magnitude). If $\sum_{\chi\in B}|\widehat{f}(\chi)|^{2}\geq(1-\mu)$, with $\mu\leq\frac{1}{8\times 2^{\varphi(\Delta)}s^{\varphi(\Delta)}}$.
One important corollary of Theorem 1.3 is that for any $s$-sparse Boolean function $f:\mathbb{Z}_{2}^{n}\to\{-1,+1\}$ the non-zero Fourier coefficients has magnitude at least $1/2^{k}$, where $k=\lceil\log_{2}s\rceil+1$. Unfortunately, such a simple corollary cannot be claimed for $s$-sparse functions $f:\mathbb{Z}_{p^{m}}^{n}\to\{-1,+1\}$ if $p^{m}\neq 2$. The main reason is that the definition of granularity is for complex numbers, rather than real numbers and hence such a lower bound cannot be directly deduced. However, borrowing results from algebraic number theory, we can obtain a lower bound on the Fourier coefficients of Boolean-valued functions from $\mathcal{G}$ to $\{-1,+1\}$, for any finite Abelian group $\mathcal{G}$.
@@
Let $f:\mathcal{G}\to\{-1,+1\}$, with Fourier sparsity $s_{f}$. Then, for any $\chi\in\operatorname{supp}(f)$, we have $|\hat{f}(\chi)|\geq\frac{1}{2^{\varphi(\Delta)/2}s^{\varphi(\Delta)/2}}$.
Remark.
One can also prove a weaker lower bound of the form $\frac{1}{2^{s_{f}/2}s_{f}^{s_{f}/2}}$, which is $\Delta$-independent; for details, see Theorem 4.5.
Observe that the lower bound in Theorem 1.5 is much lower than $1/s_{f}$. One may wonder how tight our result is. It is known that $1/s$ is tight for the case when $p=2$. For example, consider the function $AND:\mathbb{Z}_{2}^{n}\to\{-1,+1\}$. Its non-empty Fourier coefficients are either $\frac{1}{2^{n-1}}$ or $\frac{1}{2^{n-1}}$, while the empty (constant) coefficient being $1-\frac{1}{2^{n-1}}$.
@@
We prove a generalized version of the lower bound result (Theorem 1.5) for Boolean-valued functions over $\mathcal{G}$.
Finally, we design efficient algorithms for testing whether a function $f:\mathcal{G}\to\{-1,+1\}$ is $s$-sparse or “far” from $s$-sparse-Boolean.
To state our results we need to define what we mean by a function $f:\mathcal{G}\to\{-1,+1\}$ is $\epsilon$-far from $s$-sparse.
@@
A function $f:\mathcal{G}\to\{-1,+1\}$ is $\epsilon$-far from $s$-sparse-Boolean if for every $s$-sparse function $g:\mathcal{G}\to\{-1,+1\}$ the $\ell_{2}$-distance of $f$ and $g$ is at least $\sqrt{\epsilon}$. [^5] In property testing usually the distance measure used is Hamming distance between two Boolean functions. But since we are using $\ell_{2}$ distance in our other theorem, so for ease of presentation, we have defined the farness in terms of $\ell_{2}$ instead of Hamming distance. Also, note that for a pair of Boolean-valued functions the square of the $\ell_{2}$ distance and Hamming distance are the same up to a multiplicative factor of 44, see Lemma 2.43.
We say that an algorithm (property tester) $\mathcal{A}$  $\epsilon$-tests $\mathcal{C}$, for a class of functions $f:\mathcal{G}\to\{-1,+1\}$, if given access to the truth table of a function $f$, whether $f\in\mathcal{C}$, or $f$, is $\epsilon$-far from $\mathcal{C}$” can be tested using $\mathcal{A}$ with success probability (called the confidence) $\geq 2/3$. The number of queries to the truth-table of $f$ made by $\mathcal{A}$ is called the query complexity of $\mathcal{A}$.
Using the structure theorems (Theorem 1.3 and Theorem 1.4), we prove the following theorem which tests sparsity of a Boolean-valued function $f:\mathcal{G}\to\{-1,+1\}$.